OpenStudy (anonymous):

2y=x+4, what is the slope

OpenStudy (angela210793):

y=x/2+2 slope =1/2

OpenStudy (anonymous):

gradient is 1/2

OpenStudy (anonymous):

thanks i just solved it and have the same answer thanks